In 120 throws of single dice, the following distribution of faces was obtained.
Faces 1 2 3 4 5 6
Frequency 30 25 18 10 22 15
Can you say that the dice is defective?

Answer by Boreal(15235)

Ho=frequencies equal
Ha=they are not
alpha=0.05 (could choose another)
Expect the frequencies of 20 for each.
Test statistic is a Chi-Square with 5 df
if one uses the 5% level of significance, the critical value is 11.07
the test statistic is the sum of (O-E)^2/E, where E is expected and O observed.
This is
100/20=5
25/20=1.25
4/20=0.2
100/20=5
2/20=0.1
25/20=1.25
The sum is 12.8, which has a p-value between 0.01 and 0.05
conclude that the die is biased or defective in this instance.